### v1

# Online Feature Updates Improve Online (Generalized) Label Shift Adaptation

Abstract
This paper addresses the prevalent issue of label shift in an online setting with missing labels, where data distributions change over time and obtaining timely labels is challenging. While existing methods primarily focus on adjusting or updating the final layer of a pre-trained classifier, we explore the untapped potential of enhancing feature representations using unlabeled data at test-time. Our novel method, Online Label Shift adaptation with Online Feature Updates (OLS-OFU), leverages self-supervised learning to refine the feature extraction process, thereby improving the prediction model. Theoretical analyses confirm that OLS-OFU reduces algorithmic regret by capitalizing on self-supervised learning for feature refinement. Empirical studies on various datasets, under both online label shift and generalized label shift conditions, underscore the effectiveness and robustness of OLS-OFU, especially in cases of domain shifts.
1 Introduction
The effectiveness of most supervised learning models relies on a key assumption that the training data and test data share the same distribution. However, this assumption rarely holds in real-world scenarios, leading to the phenomenon of distribution shift (Quiñonero-Candela et al., 2009; Amodei et al., 2016). Previous research has primarily focused on understanding distribution shifts in offline or batch settings, where a single shift occurs between the training and test distributions (Lin et al., 2002; Shimodaira, 2000; Zadrozny, 2004; Zhang et al., 2013; Lipton et al., 2018). In contrast, real-world applications often involve test data arriving in an online fashion, and the distribution shift can continuously evolve over time. Additionally, there is another challenging issue of missing and delayed feedback labels, stemming from the online setup, where gathering labels for the streaming data in a timely manner becomes a challenging task.
To tackle the distribution shift problem, prior work often relies on additional assumptions regarding the nature of the shift, such as label shift or covariate shift (Schölkopf et al., 2012). In this paper, we focus on the common (generalized) label shift problem in an online setting with missing labels (Wu et al., 2021) . Specifically, the learner is given a fixed set of labeled training data $D_{0}\sim\mathcal{P}^{\rm train}$ in advance and trains a model $f_{0}$. During test-time, only a small batch of unlabelled test data $S_{t}\sim\mathcal{P}^{\rm test}_{t}$ arrives in an online fashion ($t=1,2,\cdots$). For the online label shift, we assume the label distribution $\mathcal{P}^{\rm test}_{t}(y)$ may change over time $t$ while the conditional distribution remains the same, i.e. $\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y)$. For example, employing MRI image classifiers for concussion detection becomes challenging as label shifts emerge from seasonal variations in the image distribution. A classifier trained during skiing season may perform poorly when tested later, given the continuous change in image distribution between skiing and non-skiing seasons. In contrast to label shift, the generalized label shift relaxes the assumption of an unchanged conditional distribution on $x$ given $y$. Instead, it assumes that there exists a transformation $h$ of the covariate, such that the conditional distribution $\mathcal{P}^{\rm test}_{t}(h(x)|y)=\mathcal{P}^{\rm train}(h(x)|y)$ stays the same. Reiterating our example, consider an MRI image classifier that undergoes training and testing at different clinics, each equipped with MRI machines of varying hardware and software versions. As a result, the images may display disparities in brightness, resolution, and other characteristics. However, a feature extractor $h$ exists, capable of mapping these variations to the same point in the transformed feature space, such that the conditional distribution $\mathcal{P}(h(x)|y)$ remains the same. In both settings, the goal of the learner is to adapt to the (generalized) label shift within the non-stationary environment, continually adjusting the model’s predictions in real-time.
Figure 1: Overview of online distribution shift adaptation. We further assume $\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y)$ for online label shift and assume the existence of an unknown feature mapping $h$ such that $\mathcal{P}^{\rm test}_{t}(h(x)|y)=\mathcal{P}^{\rm train}(h(x)|y)$ for online generalized label shift.
Existing algorithms for online label shift adaptation (OLS) primarily adopt one of two approaches: either directly reweighting of the pretrained classifier $f_{0}$, or re-training only the final linear layer of $f_{0}$ — while keeping the feature extractor frozen. Recent work (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022) have demonstrated the potential for improving feature extractors, even during test-time and in the absence of labeled data. We hypothesize that a similar effect can be harnessed in the context of (generalized) label shift, leading to the idea of improving feature representation learning during testing. In online label shift, updating the feature extractor offers two potential advantages. First, it utilizes the additional unlabeled samples, hence enhancing the sample efficiency of the feature extractor. Second, it enables the feature extractor to adapt to label shift, which is crucial to the learning process as the optimal feature extractor may depend on the underlying label distribution. Particularly in generalized label shift scenarios, where the feature transformation $h$ is often unknown, the integration of extra unlabeled test samples facilitates the learning of $h$.
Building upon this insight, this paper introduces the Online Label Shift adaptation with Online Feature Updates (OLS-OFU) framework, aimed at enhancing feature representation learning in the context of online label shift adaptation. Specifically, each instantiation of OLS-OFU incorporates a self-supervised learning method associated with a loss function denoted as $l_{\rm ssl}$ for feature representation learning, and an online label shift adaptation (OLS) algorithm to effectively address distribution shift. However, it is worth pointing out that naively integrating feature extractor updates into traditional OLS algorithms will violate the underlying key assumptions of OLS algorithms (as we show in Section 3.1), hence compromising both performance and theoretical guarantees. To address this, our OLS-OFU is carefully designed to facilitate online feature updates within the OLS paradigm. At each time step, OLS-OFU executes a modified OLS algorithm, updates the feature extractor through self-supervised learning, and subsequently refines the final linear layer. This approach enables seamless online feature updates while preserving the theoretical guarantees inherent to the original OLS algorithms.
In addition to its ease of implementation and seamless integration with existing OLS algorithms, OLS-OFU also shows strong theoretical guarantee and empirical performance. Theoretically, we demonstrate that OLS-OFU effectively reduces the loss of the overall algorithm by leveraging self-supervised learning techniques to enhance the feature extractor, thereby improving predictions for test samples at each time step $t$. Empirical evaluations on both CIFAR-10 and CIFAR-10C datasets, considering both online label shift and online generalized label shift scenarios, validate the effectiveness of OLS-OFU. It consistently outperforms its counterpart OLS methods across various datasets and distribution shift settings, underscoring its robustness across different scenarios.
2 Problem Setting & Related Work
We start with some basic notations. Let $\Delta^{K-1}$ be the probability simplex. Let $f:\mathcal{X}\to\Delta^{K-1}$ denote a classifier. Given an input $x$ from domain $\mathcal{X}$, $f(x)$ outputs a probabilistic prediction over $K$ classes. For example, $f$ can be the output from the softmax operation after any neural network. If we reweight a model $f$ by a vector $p\in\mathbb{R}^{K}$, we refer to this model as $g(\cdot;f,p)$ with $g$ denotes the method of reweighting. For any two vectors $p$ and $q$, $p/q$ denotes the element-wise division.
Online distribution shift adaptation. The effectiveness of any machine learning model $f$ relies on a common assumption that the train data $D_{0}$ and test data $D_{\text{test}}$ are sampled from the same distribution, i.e., $\mathcal{P}^{\rm train}=\mathcal{P}^{\rm test}$. However, this assumption is often violated in practice, which leads to distribution shift. This can be caused by various factors, such as data collection bias and changes in the data generation process. Moreover, once a well-trained model $f_{0}$ is deployed in the real world, it moves into the testing phase, which is composed of a sequence of time periods or time steps. The test distribution at time step $t$, $\mathcal{P}_{t}^{\rm test}$, from which test data $x_{t}$ is sampled, may vary over time. One example is that an MRI image classifier might be trained on MRI images collected during skiing season (which may have a high frequency of head concussions) but tested afterward (when the frequency of concussion is lower). The test stage can last several months until the next classifier is trained. During this test period, the distribution of MRI images may undergo continuous changes, transitioning between non-skiing and skiing seasons.
As the test-time distribution changes over time, the challenge lies in how to adjust the model continuously from $f_{t-1}$ to $f_{t}$ in an online fashion to adapt to the current distribution $\mathcal{P}^{\rm test}_{t}$. We call this problem online distribution shift adaptation and illustrate it in Figure 1. Given a total of $T$ steps in the online test stage, we define the average loss for any online algorithm $\mathcal{A}$ through the loss of the sequence of models $f_{t}$, $t\in[T]$ that are produced from $\mathcal{A}$, i.e.,
$L(\mathcal{A};\mathcal{P}_{1}^{\rm test},\cdots,\mathcal{P}_{T}^{\rm test})=\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t};\mathcal{P}_{t}^{\rm test}),$ (1)
where $\ell(f;\mathcal{P})=\mathbb{E}_{(x,y)\sim\mathcal{P}}\ell_{\rm sup}(f(x),y)$ and $\ell_{\rm sup}$ is the loss function, for example, 00-$1$ loss or cross-entropy loss for classification tasks.
If we have knowledge of $\mathcal{P}^{\rm test}_{t}$ or have enough samples from $\mathcal{P}^{\rm test}_{t}$, then the problem reduces to offline distribution shift for each single time step $t$. Alternatively, if we have a few labeled samples, the problem can be treated as classical online learning, and we can run stochastic gradient descent at the end of every time step $t$ to meet certain theoretical guarantees. However, dealing with very few unlabeled samples or even a single sample $x_{t}$ is more realistic yet technically highly non-trivial. An effective algorithm under this scenario has to utilize information from all historical data (train data $D_{0}$, validation data $D_{0}^{\prime}$ and test data up to time $t$), as well as previously deployed models $f_{0,1,\cdots,t-1}$. In this paper, we consider the challenging scenario where at each time step $t$, only a small batch of unlabeled samples $S_{t}=\{x_{t}^{1},\cdots,x_{t}^{B}\}$ is received. We formalize the algorithm $\mathcal{A}$ as: $\forall t\in[T]$,
$f_{t}:=\mathcal{A}\left(\{S_{1},\cdots,S_{t-1}\},\{f_{1},\cdots,f_{t-1}\},D_{0},D_{0}^{\prime}\right).$ (2)
In contrast to the classical online learning setup, this scenario presents a significant challenge as classical online learning literature usually relies on having access to either full or partial knowledge of the loss at each time step, i.e., $\ell_{\rm sup}(f(x_{t}),y_{t})$. In this setting, however, only a batch of unlabeled samples is provided at each time step. This lack of access to label information and loss values presents a significant challenge in accurately estimating the true loss defined in Equation 1.
Related work have studied the online distribution shift adaptation under various distribution shift assumptions, such as label shift (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023) and covariate shift (Zhang et al., 2023). In this paper, we focus on online (generlized) label shift adaptation.
Online label shift adaptation. Online label shift assumes
$\forall t\in[T],\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y),$
while the marginal distribution of the label $\mathcal{P}^{\rm test}_{t}(y)$ changes over time. This assumption is most typical when the label $y$ is the causal variable and and the feature $x$ is the observation  (Schölkopf et al., 2012). The aforementioned example of concussion detection from MRI images fits this scenario, where the presence or absence of a concussion (label) causes the observed MRI image features. Most previous methods tackle this problem through a non-trivial reduction to the classical online learning problem. Consequently, most of the online label shift algorithms (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023) study the theoretical guarantee of the algorithm via the convergence of the regret function, which could be either static regret or dynamic regret. In previous studies, the hypothesis class $\mathcal{F}$ of the prediction function $f$ is typically chosen in one of two ways. The first approach involves defining $\mathcal{F}$ as a family of post-hoc reweightings of $f_{0}$, with the parameter space comprising potential reweight vectors. Examples within this category include ROGD (Wu et al., 2021), FTH (Wu et al., 2021), and FLHFTL (Baby et al., 2023). The second approach defines $\mathcal{F}$ as a family of functions that share the same parameters in $f_{0}$ except the last linear layer, such as UOGD (Bai et al., 2022) and ATLAS (Bai et al., 2022).
Online generalized label shift adaptation. In the context of MRI image classification, where head MRI images serve as the feature $x$, variations in software or hardware across different clinics’ MRI machines can introduce discrepancies in image characteristics like brightness, contrast, and resolution. In such scenarios, the conditional probability distribution $\mathcal{P}(x|y)$ is no longer invariant. However, when a feature extractor $h$ is robust enough, it can map images into a feature space where the images from different machines have the same distributions $\mathcal{P}(h(x)|y)$ in the transformed feature space. The concept of generalized label shift, as introduced in Tachet des Combes et al. (2020), formalizes this by postulating the existence of an unknown function $h$, such that $\mathcal{P}(h(x)|y)$ remains invariant. The primary challenge in this context is to find the underlying transformation $h$. Building upon this, online generalized label shift assumes that, there exists an unknown function $h$ such that
$\forall t\in[T],\mathcal{P}^{\rm test}_{t}(h(x)|y)=\mathcal{P}^{\rm train}(h(x)|y).$
Self-supervised learning. Inspired by the extensive body of work in semi-supervised learning (Grandvalet & Bengio, 2004b; Lee et al., 2013; Laine & Aila, 2016; Gidaris et al., 2018; Miyato et al., 2018) and unsupervised representation learning (Chen et al., 2020a; He et al., 2020; Grill et al., 2020; He et al., 2022), self-supervised learning (SSL) techniques have emerged as promising tools for improving feature extraction from unlabeled data, such as applications in image classification. Likewise, when dealing with online label shift adaptation, it is crucial to leverage the unlabeled test samples $S_{1}\cup\cdots\cup S_{t-1}$ obtained from previous time steps. Ideally, these unlabeled samples could improve the feature representation learning and ultimately lead to better prediction for the test samples $S_{t}$ at time step $t$.
In this paper, we narrow our focus on three particular SSL techniques in the evaluation for classification tasks: rotation degree prediction (Gidaris et al., 2018; Sun et al., 2020), entropy minimization (Grandvalet & Bengio, 2004b; Wang et al., 2020) and MoCo (He et al., 2020; Chen et al., 2020b, 2021). It is important to note that this concept extends beyond these three SSL techniques, and the incorporation of more advanced SSL techniques has the potential to further elevate the performance. Specifically, rotation degree prediction involves initially rotating a given image by a specific degree from the set $\{0,90,180,270\}$ and the classifier is required to determine the degree by which the image has been rotated. Entropy minimization utilizes a minimum entropy regularizer, with the motivation that unlabeled examples are mostly beneficial when classes have a small overlap. MoCo is a more advanced representation learning technique, using a query and momentum encoder to learn representations from unlabeled data by maximizing the similarity of positive pairs and minimizing the similarity of negative pairs. More details of these SSL techniques are given in Appendix E.8.
3 Method
We introduce a novel online label shift adaptation algorithm OLS-OFU, that leverages self-supervised learning (SSL) to improve representation learning, while maintaining theoretical guarantees. This approach is general and can be seamlessly integrated with any existing online label shift (OLS) method. We show both theoretical and empirical improvement of OLS-OFU. Moreover, we demonstrate their effectiveness in addressing the more challenging problem of online generalized label shift adaptation.
3.1 Online Label Shift Adaptation with Online Feature Updates
Main ideas.
Before formally introducing the algorithm, we would like to discuss two crucial design principles about where to place the update feature extractor through SSL, inspired by the analysis of OLS methods in the literature. These principles are essential for maintaining both the theoretical guarantees and potential empirical performance .
Principle 1: at time step $t$, the feature extractor update using samples $S_{t}$ should occur after the OLS updates.
The main idea of ROGD, UOGD, and ATLAS is to construct an unbiased estimator for the gradient $\nabla_{f}\ell(f_{t};\mathcal{P})$ using a batch of samples $S_{t}\sim\mathcal{P}_{t}^{\rm test}$. The estimator has the form of
$\sum_{y\in\mathcal{Y}}s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y),$ (3)
where the randomness of $s_{t}$ depends on samples $S_{t}$ as constructed and is an unbiased estimator for the label marginal distribution $q_{t}$. The following proposition introduces a necessary condition of $f_{t}$ to ensure the unbiased property of the above estimator and the proof is given at Appendix B.
Proposition 1.
If $f_{t}$ is not independent of the samples $S_{t}$, the gradient estimator in Equation 3 is not guaranteed to be an unbiased estimator of the gradient $\nabla_{f}\ell(f_{t};\mathcal{P})$.
From the proposition, it is necessary to ensure $f_{t}$ is independent of $S_{t}$. Thus, the update of the feature extractor in $f_{t}$ (using SSL) using samples $S_{t}$ should occur after the online label shift (OLS) method is updated with $S_{t}$.
Principle 2: the last linear layer should be re-trained using the train set $D_{0}$ after the feature extractor update.
The key observation here is that a label-shifted distribution $\mathcal{P}^{\rm test}_{t}(y|x)$ is an exact post-hoc reweight of $\mathcal{P}^{\rm train}(y|x)$:
$\mathcal{P}^{\rm test}_{t}(y|x)=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\mathcal{P}^{\rm test}_{t}(x|y)=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\mathcal{P}^{\rm train}(x|y)$
$=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\frac{\mathcal{P}^{\rm train}(x)}{\mathcal{P}^{\rm train}(y)}\mathcal{P}^{\rm train}(y|x)\propto\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm train}(y)}\mathcal{P}^{\rm train}(y|x).$
Another observation is that $f_{0}$ trained from $D_{0}$ could be a good approximation of $\mathcal{P}^{\rm train}(y|x)$. These two observations motivate approaches like ROGD, FTH, and FLHFTL to opt for the hypothesis $f$ to be a post-hoc reweight of the pre-trained model $f_{0}$. In the context of our framework, these methods need to be adapted slightly. Instead of reweighting the original model $f_{0}$ directly, they should reweight a model within the updated feature extractor. This reweighted model should still approximate the original train data distribution $\mathcal{P}^{\rm train}(y|x)$, which can be achieved by re-training the last linear layer using the initial train data $D_{0}$.
Algorithm details.
The two principles together motivate our final algorithm: Online Label Shift adaptation with Online Feature Updates (OLS-OFU; Algorithm 1), which requires a self-supervised learning loss $\ell_{\rm ssl}$ and an online label shift adaptation algorithm (OLS) that either reweights the offline pre-trained model $f_{0}$ or updates the last linear layer [^1] As pointed out in Section 2, this is relatively general as most previous online label shift algorithms belong to these two categories.. In the training stage, we train $f_{0}$ by minimizing the supervised and self-supervised loss together defined on train data. In the test stage, OLS-OFU comprises three steps at each time step $t$: (1) running the refined version of OLS, which we refer to as OLS-R, (2) updating the feature extractor, and (3) re-training the last linear layer. The details of these three steps are elaborated below.
Algorithm 1 Online label shift adaptation with online feature updates (OLS-OFU).
  Require: An online label shift adaptation algorithm OLS, a self-supervised learning loss $\ell_{\rm ssl}$. A pre-trained model $f_{0}$ and initialize $f_{1}=f_{1}^{\prime\prime}:=f_{0}$.
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, train set $D_{0}$ and validation set $D_{0}^{\prime}$.
     1. Run the revised version of OLS, that is, OLS-R, and get $f_{t+1}^{\prime}$. (Algorithm 2 is the revised FLHFTL.)
     2. Update the feature extractor $\theta_{t}^{\rm feat}$ in $f_{t+1}^{\prime}$ by Equation 4. Replace $\theta_{t}^{\rm feat}$ in $f_{t+1}^{\prime}$ by $\theta_{t+1}^{\rm feat}$.
     3. Re-train the last linear layer by Equation 5. Calibrate the model $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear}_{t+1})$ by temperature calibration on $D_{0}^{\prime}$ and get the calibrated model $f_{t+1}^{\prime\prime}$. Output at time $t$: For the reweighting OLS methods, denote the latest reweighting vector from Step 1 is $p_{t+1}$ and we define $f_{t+1}:=g(\cdot;f_{t+1}^{\prime\prime},p_{t+1})$; else, we define $f_{t+1}:=f_{t+1}^{\prime}$.
  end for
Algorithm 2 Revised FLHFTL for online feature updates (FLHFTL-R); See the original version in (Baby et al., 2023).
  Require: Online regression oracle ALG.
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, intermediate models $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$, the validation set $D_{0}^{\prime}$, the train label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
[…]
     2. Compute $\tilde{q}_{t+1}:={\rm ALG}(s_{1},\cdots,s_{t})$
[…]
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
(1) Running the Revised OLS. At the beginning of the time $t$, we use $f_{t}^{\prime\prime}$ to denote the model within the latest feature extractor and the re-trained linear model (using data $\{S_{0},\cdots,S_{t-1}\}$). The high level idea of our framework centers on substituting the pre-trained model $f_{0}$ used in existing OLS methods with our updated model $f_{t}^{\prime\prime}$, which we call Revised OLS. To provide an overview, we examine common OLS methods (FLHFTL, FTH, ROGD, UOGD, and ATLAS) and identify two primary usecases of $f_{0}$. Firstly, all OLS methods rely on an unbiased estimator $s_{t}$ of the label distribution $q_{t}$ and $f_{0}$ is a part of the estimator. Secondly, the hypothesis $f$ is either reweighted based on $f_{0}$ or only the last linear layer in $f_{0}$ is updated. To illustrate the revised OLS formally, Algorithm  2 showcases the specific modifications applied to FLHFTL as an example. Further details regarding the revisions for ROGD, FTH, UOGD, and ATLAS can be found in Appendix C. We use $f_{t+1}^{\prime}$ to denote the model after running the revised OLS.
(2) Updating the Feature Extractor. We now introduce how to utilize the SSL loss $\ell_{\rm ssl}$ to update the feature extractor based on any incoming batch of unlabeled test data $S_{t}$ at timestep $t$. Specifically, let $\theta^{\rm feat}_{t}$ denote the parameters of the feature extractor in $f_{t+1}^{\prime}$. The update of $\theta^{\rm feat}_{t+1}$ at time $t$ is given by:
$\theta_{t+1}^{\rm feat}:=\theta_{t}^{\rm feat}-\eta\cdot\nabla_{\theta^{\rm feat}}\ell_{\rm ssl}(S_{t};f_{t+1}^{\prime}).$ (4)
Then we update the feature extractor in $f_{t+1}^{\prime}$ with $\theta_{t+1}^{\rm feat}$.
(3) Re-training Last Linear Layer. With the updated feature extractor $\theta_{t+1}^{\rm feat}$, it is important to re-train the last linear layer $\theta^{\rm linear}_{t+1}$ to align with the new feature extractor. The re-training starts from random initialization, while the feature extractor remains frozen. The training objective of $\theta^{\rm linear}_{t+1}$ is to minimize the average cross-entropy loss over train data $D_{0}$. We denote the model with the frozen feature extractor $\theta_{t+1}^{\rm feat}$ as $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear})$. The objective for re-training last linear layer can be written as follows:
$\theta^{\rm linear}_{t+1}:=\arg\min_{\theta_{\rm linear}}\sum_{(x,y)\in D_{0}}\ell_{\rm ce}\bigg{(}f(x|\theta^{\rm feat}_{t+1},\theta^{\rm linear}),y\bigg{)}.$ (5)
We calibrate the model $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear}_{t+1})$ by temperature calibration (Guo et al., 2017) using the validation set $D_{0}^{\prime}$ and denote the model after calibration as $f_{t+1}^{\prime\prime}$. This re-training step is essential to ensure the model is a calibrated model of estimating $\mathcal{P}^{\rm train}(y|x)$ rather than other distributions.
Ouput at time $t$. Finally, we define $f_{t+1}$ for the next time step prediction. For the reweighting OLS methods (ROGD, FTH, FLHFTL), denote the latest reweighting vector from Step 1 is $p_{t+1}$ and define $f_{t+1}:=g(\cdot;f_{t+1}^{\prime\prime},p_{t+1})$. For those which optimize the last linear layer (UOGD, ATLAS), we define $f_{t+1}:=f_{t+1}^{\prime}$ – as stated in Principle 2 in Section 3.1, the $f_{t}^{\prime\prime}$ including retrained last linear layer serves specially for the reweighting OLS method.
Batch accumulation for SSL that needs large batches. The loss of SSL such as contrastive learning (Chen et al., 2020a; He et al., 2020) is defined from a batch of data and the effective batch is necessary to be large. We introduce a simple variant of Algorithm 1: we process Step 2-3 every BA time steps by the samples collected in the period.
3.2 Performance Guarantee
The original OLS methods exhibit theoretical guarantees in terms of regret convergence for online label shift setting, where $\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y)$ is invariant. With the incorporation of the additional online feature update step, OLS-OFU demonstrates analogous theoretical results. By comparing the theoretical results between OLS and OLS-OFU, we can gain insights into potential enhancements from OLS to OLS-OFU. Due to limited space, we provide the theoretical results pertaining to FLHFTL-OFU here, and present the theoretical results for ROGD-OFU, FTH-OFU, UOGD-OFU, and ATLAS-OFU in Appendix D.
Theorem 2.
[Regret convergence for FLHFTL-OFU] Suppose we choose the OLS subroutine in Algorithm 2 to be FLHFTL from Baby et al. (2023). Let $f_{t}^{\rm flhftl-ofu}$ be the output at time step $t-1$ from Algorithm 1, that is $g(\cdot;f_{t}^{\prime\prime},\frac{\tilde{q}_{t}}{q_{0}})$. Let $\sigma$ be the smallest among the the minimum singular values of invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$. Then under Assumptions 1 and 2 in Baby et al. (2023), FLHFTL-OFU has the guarantee below: $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})-\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})\right]$ $\leq O\left(\frac{K^{1/6}V_{T}^{1/3}}{\sigma^{2/3}T^{1/3}}+\frac{K}{\sigma\sqrt{T}}\right),$ (6)
where $V_{T}:=\sum_{t=1}^{T}\|q_{t}-q_{t-1}\|_{1}$, $K$ is the number of classes, and the expectation is taken w.r.t. randomness in the revealed co-variates. This result is attained without prior knowledge of $V_{T}$.
To ease comparison, we state the guarantee for the original OLS method FLHFTL (Baby et al., 2023):
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl};\mathcal{P}_{t}^{\rm test})\right]-\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{0},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})$ $\leq O\left(\frac{K^{1/6}V_{T}^{1/3}}{\sigma^{2/3}T^{1/3}}+\frac{K}{\sigma\sqrt{T}}\right),$ (7)
Recall that the objective function for the online label shift problem is defined as the average loss in Equation 1. Both theorems establish the convergence in terms of this average loss. In the event that $f_{t}^{\prime\prime}$ ($t\in[T]$) from the online feature updates yield improvements:
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})\right]<\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{0},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test}),$ (8)
then it guarantees that the loss of FLHFTL-OFU will converge to a smaller value, resulting in enhanced performance compared to FLHFTL. We substantiate this improvement through empirical evaluation in Section 4. For other OLS methods such as ROGD-OFU, FTH-OFU, UOGD-OFU, and ATLAS-OFU, a similar analysis can be derived and we present them in Appendix D.
3.3 Online Feature Updates Improve Online Generalized Label Shift Adaptation
When we have knowledge of the feature map $h$ under generalized label shift, the problem simplifies to the classical online label shift scenario. However, the more challenging situation arises when $h$ remains unknown. Due to the violation of the label shift assumption, standard OLS methods can perform arbitrarily bad. Fortunately, existing research in test-time training (TTT) (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022) demonstrates that feature updates driven by SSL align the source and target domains in feature space. When the source and target domains achieve perfect alignment, such feature extractor effectively serves as the feature map $h$ as assumed in generalized label shift. Therefore, the sequence of feature extractors in $f_{1},\cdots,f_{T}$ generated by Algorithm 1 progressively approximates the underlying $h$. This suggests that, compared to the original OLS, OLS with online feature updates (Algorithm 1) experiences a milder violation of the label shift assumption within the feature space and is actually expected to have better performance in the setting of online generalized label shift. Indeed, as demonstrated in Section 4, we can observe significant improvements in OLS-OFU compared to standard OLS for online generalized label shift.
4 Experiment
Figure 2: Evaluation of OLS and OLS-OFU.
In this section, we empirically evaluate how OLS-OFU improves the original OLS methods on both online label shift and online generalized label shift [^2] Code released at https://anonymous.4open.science/r/online_label_shift_with_online_feature_updates-3A1F/.. The experiment is performed on CIFAR-10 (Krizhevsky et al., 2009), STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), EuroSAT (Helber et al., 2019), and CIFAR-10C (Hendrycks & Dietterich, 2019). We systematically vary the distribution shift processes and self-supervised learning (SSL) techniques to assess the effectiveness of the proposed method.
4.1 Experiment Set-up
Dataset and Label Shift Settings. For online label shift, we evaluate the efficacy of our algorithm on CIFAR-10 (Krizhevsky et al., 2009), STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), and EuroSAT (Helber et al., 2019), which have 10 categories of images. For each dataset, we split the original train set into the offline train (i.e., $D_{0}$) and validation sets (i.e., $D^{\prime}_{0}$) following a ratio of $4:1$. At the online test unlabeled batches are sampled from the test set. For online generalized label shift, the offline train and validation sets are the same CIFAR-10 images. The test unlabeled batches are drawn from CIFAR-10C, a benchmark with the same objects as CIFAR-10 but with various types of corruption. We experiment with three types of corruptions with CIFAR-10C: Gaussian noise, Fog and Pixelate.
We follow Bai et al. (2022) and Baby et al. (2023) to simulate the online label distribution shift with two shift patterns: Sinusoidal shift and Bernoulli shift. Specifically, given two label distribution vectors $q$ and $q^{\prime}$, we simulate the label marginal distributions at time $t$ as a weighted combination of them: $q_{t}:=\alpha_{t}q+(1-\alpha_{t})q^{\prime}$. In Sinusoidal shift, $\alpha_{t}=\sin\frac{i\pi}{L}$ (periodic length $L=\sqrt{T}$, $i=t\textnormal{ mod }L$) while in Bernoulli shift, $\alpha_{t}$ is a random bit (either 0 or 1), where the bit switches $\alpha_{t}=1-\alpha_{t-1}$ with probability $p=1-\frac{1}{\sqrt{T}}$. In our experiments, we set the initial distribution vector $q$ and $q^{\prime}$ with $\frac{1}{K}(1,\cdots,1)$ and $(1,0,\cdots,0)$. To sample the batch test data at time $t$, we first sample a batch of labels (not revealed to the learner) according to $q_{t}$. Then given each label we can sample an image from the test set, and collect this batch of images without labels as $S_{t}$. We experiment with $T=1000$ and batch size $B=10$ at each time step, following Baby et al. (2023).
Evaluation Metric. We report the average error during test, i.e., $\frac{1}{TB}\sum_{t=1}^{T}\sum_{x_{t}\in S_{t}}\mathbbm{1}\left(f_{t}(x_{t})\neq y_{t}\right)$, where $(x_{t},y_{t})\sim\mathcal{P}_{t}^{\rm test}$, to approximate $\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t};\mathcal{P}^{\rm test}_{t})$ for the evaluation efficiency. This is valid for large $T$ due to its exponential concentration rate by the Azuma–Hoeffding inequality.
Baselines and OLS Methods. We perform an extensive evaluation of 6 OLS methods in the literature: FTFWH, FTH, ROGD, UOGD, ATLAS, and FLHFTL. We further report the performance of our method OLS-OFU (Algorithm 1) applied on top of each OLS. OLS-OFU is instantiated with 3 common SSL methods introduced in Section 2: rotation degree prediction, entropy minimization, and MoCo. We run the batch accumulation variant of Algorithm 1 for MoCo and set BA=$50$; please see the ablation for the selection of BA in Appendix E.8.1. Additionally, we report two baseline scores, Base and OFU. The Base uses the fixed pre-trained model $f_{0}$ to predict the labels at all time steps. The second is online feature updates (OFU) only, where at time step $t$ we only update the features (Step 2 in Algorithm 1) without utilizing OLS algorithms.
4.2 Results
Main Results: comparison between OLS-OFU and baselines under online (generalized) label shift. Figure 2(a) shows the performance comparison between OLS-OFU, implemented with three SSL methods, and their corresponding OLS counterparts on CIFAR-10 under the scenario of classical online label shift. Figure 2(b) shows the results on three more datasets with SSL technique in OLS-OFU being rotation degree prediction. The online shift pattern in the setting of Figure 2(a-b) is sinusoidal shift. From Figure 2(a-b), we can observe that OLS-OFU outperforms all other baselines. Notably, OLS-OFU, when integrated with three SSL methods, consistently outperforms their counterpart OLS method across all of the OLS methods. This emphasizes the importance of improving representation learning within label shift adaptation and underscores the robustness of our framework OLS-OFU. Results for Bernoulli shift in Appendix E.3 show similar conclusion.
Furthermore, we report the performance of OLS-OFU in the context of online generalized label shift. In this scenario, the test images exhibit three types of domain shifts in CIFAR-10C — Gaussian noise, Fog, and Pixelation—with mild severity. We present the result under Sinusoidal shift with the SSL method used in OLS-OFU as rotation degree prediction. Similar to the scenario of online label shift, the superior performance of OLS-OFU compared with Base, OFU, and OLS demonstrates its superiority (Figure 2(c)). Particularly, OLS-OFU consistently outperforms the original OLS methods by an even larger gap than the one occurring at the classic label shift setting. We hypothesize that the inclusion of SSL methods within OLS-OFU aids the learning of the unknown mapping $h$ inherent in the generalized label shift assumption. As evaluated in Sun et al. (2020), the label shift assumption in the feature space of $\theta^{\rm feat}_{t}$ is violated milder than the one of $\theta^{\rm feat}_{0}$ — the more severe violation of $\theta^{\rm feat}_{0}$ explains why some OLS methods perform even worse than OFU. Results for other SSL methods and online shift patterns show a similar pattern in Appendix E.4.
Moreover, as an ablation study of OLS-OFU, the comparison between OLS-OFU and either OLS or OFU individually underscores the specific benefits of both the OFU and OLS modules in addressing the challenges posed by the online (generalized) label shift problem.
Figure 3: Empirical examination for the holdness of Equation 8. Clean denotes the experiment on CIFAR-10 and others denotes the corruption type. They are paired with two online shift patterns: Sinusoidal and Bernoulli.
Empirical validation of Equation 8. In Section 3.2, we argued that when the inequality in Equation 8 holds, the loss of FLHFTL-OFU exhibits a tighter upper bound compared to FLHFTL. Figure 3 presents the RHS (corresponds to OLS) and LHS (corresponds to OLS-OFU with SSL loss as rotation degree prediction) of Equation 8. We perform the study over eight different settings and varies the domain shift and online shift patterns. It is evident that OLS-OFU yields improvements on the baseline of the regret as shown in Equation 8. Appendix E.6 validates this inequality for other SSL techniques. We further examine how the improvement from Equation 8 (RHS - LHS) is correlated the improvement from OLS to OLS-OFU. For each OLS method, we compute the Pearson coefficient (within the range of [-1, 1]) between these two terms across all experimental setups (varying dataset, online shift pattern, and SSL methods). The calculated coefficients are 0.70, 0.76, 0.54, 0.63, 0.58, and 0.75 when OLS is FTFWH, FTH, ROGD, ATLAS, UOGD, or FLHFTL, respectively. These values indicate strong positive correlations, underscoring how improvements in regret align with empirical performance improvements.
Results on CIFAR-10C with higher severity. We evaluate all methods on CIFAR-10C with higher severity of domain shifts. Figure 4 shows the results when the test images are corrupted by Gaussian noise [^3] Please refer to Appendix E.5 for results involving higher levels of severity, other corruption types, and SSL techniques.. Comparing with Figure 2(c), it is clear that as the severity increases, OLS-OFU enlarges the performance gap compared to OLS. However, it might be worth pointing out that neither OLS nor OLS-OFU are better than OFU.
Figure 4: Results on CIFAR-10C for a high level of severity.
[…]
5 Conclusion
We focus on online (generalized) label shift adaptation and present a novel framework Online Label Shift adaptation with Online Feature Updates (OLS-OFU), which harnesses the power of self-supervised learning to enhance feature representations dynamically during testing, leading to improved predictive models and better test time performance.
Discussion and future work. One promising direction is to extend the idea of this paper to online covariate shift, for example, the algorithm in Zhang et al. (2023) freezes the feature extractor and only updates the linear layer. Another possible direction is to consider a more realistic domain shift scenario within the generalized label shift setting — domain shift types may vary over time or be even more challenging, such as shifting from cartoon images to realistic images.
Acknowledgements
RW and KQW are supported by grants from LinkedIn, DARPA Geometries of Learning, and the National Science Foundation NSF (1934714). RW is also supported by grants from the National Science Foundation NSF (2241100), Army Research Office ARO MURI (W911NF2110317), and Office of Naval Research ONR (N00014-20-1-2334).
Appendix A Further Related Work
Offline distribution shift and domain shift. Offline label shift and covariate shift have been studied for many years. Some early work (Saerens et al., 2002; Lin et al., 2002) assumes the knowledge of how the distribution is shifted. Later work (Shimodaira, 2000; Zadrozny, 2004; Huang et al., 2006; Gretton et al., 2009; Lipton et al., 2018; Alexandari et al., 2020; Azizzadenesheli et al., 2019; Garg et al., 2020) relaxes this assumption and estimates this knowledge from unlabeled test data. A recent work by (Garg et al., 2023) considers a relaxed version of offline label shift problem where the class-conditionals can change between train and test domains in a restrictive way. Extending their results to the case of relaxed online version of generalized label shift is an interesting future direction.
Online distribution shift with provable guarantees. There has been several work modeling online distribution shift as the classic online learning problem (Wu et al., 2021; Bai et al., 2022; Baby et al., 2023; Zhang et al., 2023), which leverage the classical online learning algorithms (Shalev-Shwartz, 2012; Besbes et al., 2015; Baby & Wang, 2022) to bound the static or dynamic regret. However, none of them updates the feature extractor in a deep learning model but only the last linear layer or the post-hoc linear reweighting vectors. Our proposed method OLS-OFU utilizes the deep learning SSL to improve the feature extractor, which brings better performance.
Domain shift adaptation within online streaming data. When we consider the most authentic online learning setup where the learner only receives the unlabeled samples, the most representative idea is test-time training (Sun et al., 2020; Wang et al., 2020; Liu et al., 2021; Niu et al., 2022), which utilizes a (deep learning) self-supervised loss to online update the model. However, it focuses on how to adapt to a fixed domain shifted distribution from online streaming data and is not designed for how to adapt to continuous distribution changes during the test stage, while our algorithm concentrates the later problem. Besides test-time training, Hoffman et al. (2014) and (Mullapudi et al., 2019) study the online domain shift for specific visual applications.
Algorithm 3 Revised ROGD for online feature updates ROGD-R. See the original version in Equation 7 and Equation 8 in (Wu et al., 2021).
  Require: Learning rate $\eta$.
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, and intermediate model $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$ from step 3 in Algorithm 1, the validation set $D_{0}^{\prime}$, the training label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
[…]
     2. Grab the weight $p_{t}$ from $f_{t}$.
[…]
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
Algorithm 4 Revised FTH for online feature updates (FTH-R). See the original version in Equation 9 in (Wu et al., 2021).
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, and intermediate model $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$ from step 3 in Algorithm 1, the validation set $D_{0}^{\prime}$, the train label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
[…]
     2. Compute $p_{t+1}=\frac{1}{t}\sum_{\tau=1}^{t}s_{\tau}$
[…]
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
Algorithm 5 Revised UOGD for online feature updates (UOGD-R). See the original version in Equation 9 in (Bai et al., 2022).
  Require: The learning rate $\eta$.
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, and intermediate model $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$ from step 3 in Algorithm 1, the validation set $D_{0}^{\prime}$, the train label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
[…]
     2. Grab the weight $w_{t}$ from the last linear layer of $f_{t}$.
[…]
     4. Let $f_{t+1}$ be $f(\cdot|\theta^{\rm feat}_{t},w_{t+1})$.
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
Algorithm 6 Revised ATLAS for online feature updates (ATLAS-R). See the original version in Equation 9 in (Bai et al., 2022).
  Require: The learning rate pool $\mathcal{H}$ with size N; Meta learning rate $\varepsilon$; $\forall i\in[N]$, $p_{1,i}=1/N$ and $w_{1,i}=\theta^{\rm linear}_{f_{0}}$.
  for $t=1,\cdots,T$ do
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, and intermediate model $\{f_{1}^{\prime\prime},\cdots,f_{t}^{\prime\prime}\}$ from step 3 in Algorithm 1, the validation set $D_{0}^{\prime}$, the train label marginal $q_{0}:=\mathcal{P}^{\rm train}(y)$.
[…]
     for $i\in[N]$ do
[…]
     end for
     3. Update weight $p_{t+1}$ according to $p_{p_{t,i}}\propto\exp(-\varepsilon\sum_{\tau=1}^{t-1}\hat{R}_{\tau}(\mathbf{w}_{\tau,i}))$
     3. Compute $w_{t+1}=\sum_{i=1}^{N}p_{t+1,i}w_{t+1,i}$. Let $f_{t+1}$ be $f(\cdot|\theta^{\rm feat}_{t},w_{t+1})$.
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
Appendix B Proof of Proposition 1
The proof of Proposition 1.
We first write the derivation that $\sum_{y\in\mathcal{Y}}s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)$ is an unbiased estimator of $\nabla_{f}\ell(f_{t};\mathcal{P})$ when $f_{t}$ is independent of $S_{t}$.
$\mathbb{E}_{S_{t}}\left[\sum_{y\in\mathcal{Y}}s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)\right]$ $=\sum_{y\in\mathcal{Y}}\mathbb{E}_{S_{t}}\left[s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)\right]$
$=\sum_{y\in\mathcal{Y}}\mathbb{E}_{S_{t}}\left[s_{t}[y]\right]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)$
$=\sum_{y\in\mathcal{Y}}q_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)$
$=\sum_{y\in\mathcal{Y}}q_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm test}_{t}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y)$
$=\nabla_{f}\ell(f_{t};\mathcal{P}).$
The third equality is as how $s_{t}$ is constructed. The fourth equality holds by the label shift assumption. We can check the second equality: if $f_{t}$ is dependent of $S_{t}$, the correctness of the second equality is not guaranteed and so is the overall unbiased property derivation.
Appendix C The Revision for Previous Online Label Shift Adaptation Algorithms
The revised algorithms to be used in the main algorithm OLS-OFU (Algorithm 1) are FTH-R (Algorithm 4), UOGD-R (Algorithm 5), ROGD-R (Algorithm 3), ATLAS-R (Algorithm 6).
Appendix D Theorems for OLS and Proofs
In this section, we present the theoretical results of FLHFTL-OFU, ROGD-OFU, FTH-OFU, UOGD-OFU, ATLAS-OFU and their proofs. The proofs are mostly the same as the proofs for the original algorithms with small adjustments. As our results are not straight corollaries for the original theorems, we write the full proofs here for the completeness.
D.1 Theorem for FLHFTL-OFU
Before proving Theorem 2 (in Section 3.2 ) we recall the assumption from Baby et al. (2023) for convenience. We refer the reader to Baby et al. (2023) for justifications and further details of the assumptions.
Assumption 1.
Assume access to the true label marginals $q_{0}\in\Delta_{K}$ of the offline train data and the true confusion matrix $C\in\mathbb{R}^{K\times K}$. Further the minimum singular value $\sigma_{min}(C)=\Omega(1)$ is bounded away from zero.
Assumption 2 (Lipschitzness of loss functions).
Let $\mathcal{D}$ be a compact and convex domain. Let $r_{t}$ be any probabilistic classifier. Assume that $L_{t}(p):=E\left[\ell(g(\cdot;r_{t},p/q_{0})|x_{t}\right]$ is $G$ Lipschitz with $p\in\mathcal{D}\subseteq\Delta_{K}$, i.e, $L_{t}(p_{1})-L_{t}(p_{2})\leq G\|p_{1}-p_{2}\|_{2}$ for any $p_{1},p_{2}\in\mathcal{D}$. The constant $G$ need not be known ahead of time.
See 2 Proof:
The algorithm in Baby et al. (2023) requires that the estimate $s_{t}$ in Line 1 of Algorithm 2 is unbiased estimate of the label marginal $q_{t}$. Since $f_{t}^{\prime\prime}$ in Algorithm 2 is independent of the sample $S_{t}$, and since we are working under the standard label shift assumption, due to Lipton et al. (2018) we have that $C_{f_{t}^{\prime\prime},D_{0}^{\prime}}^{-1}\cdot\frac{1}{|S_{t}|}\sum_{x_{t}\in S_{t}}f_{t}^{\prime\prime}(x_{t})$ forms an unbiased estimate of $E_{x\sim\mathcal{P}_{t}^{\text{test}}}[f_{t}^{\prime\prime}(x)]$. Further, from Lipton et al. (2018), the reciprocal of standard deviation of this estimate is bounded below by minimum of the singular values of confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$.
Let $\tilde{q_{t}}$ be the estimate of the label marginal maintained by FLHFTL. By Lipschitzness, we have that
$E[\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})-\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0})]$ $=E[L_{t}(\tilde{q}_{t})]-E[L_{t}(q_{t})]$ (9)
$\leq G\cdot E[\|\tilde{q}_{t}-q_{t}\|_{2}],$ (10)
where the last line is via Assumption 2.
$\sum_{t=1}^{T}E[\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})-\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0})]$ $\leq\sum_{t=1}^{T}G\cdot E[\|\tilde{q}_{t}-q_{t}\|_{2}]$ (11)
$\leq\sum_{t=1}^{T}G\sqrt{E\|\tilde{q}_{t}-q_{t}\|_{2}^{2}}$ (12)
$\leq G\sqrt{T\sum_{t=1}^{T}E[\|\tilde{q}_{t}-q_{t}\|_{2}^{2}]}$ (13)
$=\tilde{O}\Bigg{(}\Bigg{.}K^{1/6}T^{2/3}V_{T}^{1/3}(1/\sigma^{2/3}_{min}(C))+\sqrt{KT}/\sigma_{min}(C)\Bigg{.}\Bigg{)},$ (14)
where the second line is due to Jensen’s inequality, third line by Cauchy-Schwartz and last line by Proposition 16 in Baby et al. (2023). This finishes the proof.
D.2 Theorem for ROGD-OFU
We state the assumptions first for the later theorems. These assumptions are similar to Assumption 1-3 in (Wu et al., 2021).
Assumption 3.
$\forall\mathcal{P}\in\{\mathcal{P}^{\rm train},\mathcal{P}^{\rm test}_{1},\cdots,\mathcal{P}^{\rm test}_{T}\}$, ${\rm diag}(C_{f,\mathcal{P}})$ is differentiable with respect to $f$.
Assumption 4.
$\forall t\in[T]$, $\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})$ is convex in $p$, where $f_{t}^{\prime\prime}$ is defined in Algorithm 1.
Assumption 5.
$\sup_{p\in\Delta^{K-1},i\in[K],t\in[T]}\|\nabla_{p}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\|_{2}$ is finite and bounded by $L$.
Theorem 3 (Regret convergence for ROGD-OFU).
If we run Algorithm 1 with ROGD-R (Algorithm 3) and $\eta=\sqrt{\frac{2}{T}}\frac{1}{L}$, under Assumption 3, 4, 5, ROGD-OFU satisfies the guarantee
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm ogd-ofu};\mathcal{P}_{t}^{\rm test})\right]-\min_{p\in\Delta_{K}}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\right]\leq\sqrt{\frac{2}{T}}L.$ (15)
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm ogd};\mathcal{Q}_{t})\right]-\min_{p\in\Delta_{K}}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;p,f_{0},q_{0});\mathcal{Q}_{t})\right]\leq\sqrt{\frac{2}{T}}L.$ (16)
Proof: For any fixed $p$,
$\ell(f_{t}^{\rm rogd-ofu};\mathcal{P}_{t}^{\rm test})-\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})$ $=\ell(g(\cdot;f_{t}^{\prime\prime},p_{t}/q_{0});\mathcal{P}_{t}^{\rm test})-\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})$
$\leq(p_{t}-p)\cdot\nabla_{p}\ell(g(\cdot;f_{t}^{\prime\prime},p_{t}/q_{0});\mathcal{P}_{t}^{\rm test})$
$=(p_{t}-p)\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}\mathbb{E}_{S_{t}}[s_{t}|S_{1},\cdots,S_{t-1}]$
$=\mathbb{E}_{S_{t}}[(p_{t}-p)\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}|S_{1},\cdots,S_{t-1}],$
where the last inequality holds by the fact that $(p_{t}-p)\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}$ is independent of $\{S_{1},\cdots,S_{t-1}\}$. To bound $(p_{t}-p)\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}$,
$\|p_{t+1}-p\|_{2}^{2}$ $=\|{\rm Prof}_{\Delta^{K-1}}(p_{t}-\eta\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t})-p\|_{2}^{2}$
$\leq\|p_{t}-\eta\cdot J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}-p\|_{2}^{2}$
$=\|p_{t}-p\|_{2}^{2}+\eta^{2}\|J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}\|_{2}^{2}-2\eta(p_{t}-p)\cdot(J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}).$
This implies
$(p_{t}-p)\cdot(J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t})\leq\frac{1}{2\eta}(\|p_{t}-p\|_{2}^{2}-\|p_{t+1}-p\|_{2}^{2})+\frac{\eta}{2}\|J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}\|_{2}^{2}$
Thus
$\mathbb{E}_{S_{1},\cdots,S_{T}}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm rogd-ofu};\mathcal{P}_{t}^{\rm test})-\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\right]$
$\leq\mathbb{E}_{S_{1},\cdots,S_{T}}\left[\frac{1}{T}\sum_{t=1}^{T}\frac{1}{2\eta}(\|p_{t}-p\|_{2}^{2}-\|p_{t+1}-p\|_{2}^{2})+\frac{\eta}{2}\|J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}\|_{2}^{2}\right]$
$\leq\frac{1}{2\eta T}\|p_{1}-p\|_{2}^{2}+\frac{\eta}{2T}\sum_{t=1}^{T}\mathbb{E}_{S_{1},\cdots,S_{t}}[\|J_{p,f_{t}^{\prime\prime}}(p_{t})^{\top}s_{t}\|_{2}^{2}]$
$\leq\frac{1}{\eta T}+\frac{\eta L^{2}}{2}=\sqrt{\frac{2}{T}}L.$
This bound holds for any p. Thus,
$\mathbb{E}_{S_{1},\cdots,S_{T}}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm rogd-ofu};\mathcal{P}_{t}^{\rm test})\right]-\min_{p\in\Delta^{K-1}}\mathbb{E}_{S_{1},\cdots,S_{T}}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\right]\leq\sqrt{\frac{2}{T}}L.$
D.3 Theorem for FTH-OFU
We begin with two assumptions.
Assumption 6.
For any $\mathcal{P}^{\rm test}$ s.t. $\mathcal{P}^{\rm test}(x|y)=\mathcal{P}^{\rm train}(x|y)$, denote $q_{t}:=(\mathcal{P}^{\rm test}_{t}(y=k):k\in[K])$ and then
$\|q_{t}-\arg\min_{p\in\Delta^{K-1}}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}^{\rm test})\|\leq\delta.$
Assumption 7.
$\forall\mathcal{P}^{\rm test}$ s.t. $\mathcal{P}^{\rm test}(x|y)=\mathcal{P}^{\rm train}(x|y)$, $\sup_{p}\|\nabla_{p}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}^{\rm test})\|\leq L$
Theorem 4 (Regret convergence for FTH-OFU).
If we run Algorithm 1 with FTH-R (Algorithm 4) and assume $\sigma$ is no larger than the minimum singular value of invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$, under Assumption 6 and 7 with $\delta=0$, FTH-OFU satisfies the guarantee that with probability at least $1-2KT^{-7}$ over samples $S_{1}\cup\cdots\cup S_{T}$,
$\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm fth-ofu};\mathcal{P}_{t}^{\rm test})-\min_{p\in\Delta_{K}}\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\leq O\left(\frac{\log T}{T}+\frac{1}{\sigma}\sqrt{\frac{K\log T}{T}}\right),$ (17)
where $K$ is the number of classes.
Proof: Denote $q_{t}:=(\mathcal{P}^{\rm test}_{t}(y=k):k\in[K])$. By the Hoeffding and union bound, we have
$\mathbb{P}\left(\forall t\leq T,\|p_{t+1}-\frac{1}{t}\sum_{\tau=1}^{t}q_{\tau}\|\leq\sqrt{K}\varepsilon_{t}\right)\geq 1-\sum_{t=1}^{T}2M\exp\left(-2\varepsilon_{t}^{2}t/\sigma^{2}\right).$
This implies that with probability at least $1-\sum_{t=1}^{T}2M\exp\left(-2\varepsilon_{t}^{2}t/\sigma^{2}\right)$, $\forall p$,
$\sum_{t=1}^{T}\ell(p_{t};\mathcal{P}_{t}^{\rm test})-\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})$
$\leq\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{1}{t}\sum_{\tau=1}^{t}q_{\tau}/q_{0});\mathcal{P}_{t}^{\rm test})-\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})+L\sqrt{M}\cdot\sum_{t=1}^{T}\varepsilon_{t}$
$\leq\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{1}{t-1}\sum_{\tau=1}^{t-1}q_{\tau}/q_{0});\mathcal{P}_{t}^{\rm test})-\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{1}{t}\sum_{\tau=1}^{t}q_{\tau}/q_{0});\mathcal{P}_{t}^{\rm test})+L\sqrt{M}\cdot\sum_{t=1}^{T}\varepsilon_{t}$
$\leq\sum_{t=1}^{T}L\left\|\frac{1}{t-1}\sum_{\tau=1}^{t-1}q_{\tau}-\frac{1}{t}\sum_{\tau=1}^{t}q_{\tau}\right\|+L\sqrt{M}\cdot\sum_{t=1}^{T}\varepsilon_{t}$
$\leq\sum_{t=1}^{T}\frac{L}{t}\left\|\frac{1}{t-1}\sum_{\tau=1}^{t-1}q_{\tau}-q_{t}\right\|+L\sqrt{M}\cdot\sum_{t=1}^{T}\varepsilon_{t}$
$\leq\sum_{t=1}^{T}\frac{2L}{t}+L\sqrt{M}\cdot\sum_{t=1}^{T}\varepsilon_{t}.$
If we take $\varepsilon_{t}=2\sigma\sqrt{\frac{\ln T}{T}}$, the above is equivalent to: with probability at least $1-2KT^{-7}$
$\frac{1}{T}\sum_{t=1}^{T}\ell(p_{t};\mathcal{P}_{t}^{\rm test})-\min_{p}\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},p/q_{0});\mathcal{P}_{t}^{\rm test})\leq 2L\frac{\ln T}{T}+4L\sigma\sqrt{\frac{K\ln T}{T}}$
D.4 Theorems for UOGD-OFU and ATLAS-OFU
Theorem 5.
[Regret convergence for UOGD-OFU] Let $f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w)$ denote a network with the same feature extractor as that of $f_{t}^{\prime\prime}$ and a last linear layer with weight $w$. Let $f^{\rm uogd-ofu}=f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})$, where $w_{t}$ is the weight maintained at round $t$ by Algorithm 5. If we run Algorithm 1 with UOGD in (Bai et al., 2022) and let step size be $\eta$, then under the same assumptions as Lemma 1 in (Bai et al., 2022), UOGD-OFU satisfies that
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f^{\rm uogd-ofu};\mathcal{P}^{\rm test}_{t})-\frac{1}{T}\sum_{t=1}^{T}\min_{w\in\mathcal{W}}\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}^{\rm test}_{t})\right]\leq O\left(\frac{K\eta}{\sigma^{2}}+\frac{1}{\eta T}+\sqrt{\frac{V_{T,\ell}}{T\eta}}\right),$ (18)
where $V_{T,\ell}:=\sum_{t=2}^{T}\sup_{w\in\mathcal{W}}|\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}_{t}^{\rm test})-\ell(f(\cdot;\theta^{\rm feat}_{f_{t-1}^{\prime\prime}},w);\mathcal{P}_{t-1}^{\rm test})|$, $\sigma$ denotes the minimum singular value of the invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$ and $K$ is the number of classes and the expectation is taken with respect to randomness in the revealed co-variates.
Proof Sketch: Recall that $\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}^{\rm test}_{t}):=E_{(x,y)\sim\mathcal{P}_{t}^{\rm test}}\ell_{\rm ce}\left(f(x|\theta^{\rm feat}_{f_{t}^{\prime\prime}},w),y\right)$.
This guarantee follows from the arguments in Bai et al. (2022) from two basic facts below:
  1. 1.
    The risk $\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}^{\rm test}_{t})$ is convex in $w$ over a convex and compact domain $\mathcal{W}$.
  2. 2.
    It is possible to form unbiased estimates $\hat{G}_{t}(w)\in\mathbb{R}^{K}$ such that $E[\hat{G}_{t}(w)|S_{1:t-1}]=E_{((x,y)\sim\mathcal{P}_{t}^{\rm test})}\nabla_{w}\ell_{\rm ce}\left(f(x|\theta^{\rm feat}_{f_{t}^{\prime\prime}},w),y\right)$.
Hence we proceed to verify these two facts in our setup. Fact 1 is true because the cross-entropy loss is convex in any subset of the simplex and the last linear layer weights only defines an affine transformation which preserves convexity.
For fact 2, note that the $f_{t}^{\prime\prime}$ only uses the data until round $t-1$. So by the same arguments in Bai et al. (2022), using the BBSE estimator defined from the classifier $f_{t}^{\prime\prime}$, the unbiased estimate of risk gradient can be defined.
Let $w_{t}$ be the weight of the last layer maintained by UOGD at round t. Let $u_{1:T}$ be any sequence in $\mathcal{W}$. Consequently we have for any round,
$\ell(f^{\rm uogd-ofu};\mathcal{P}^{\rm test}_{t})-\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},u_{t}))$ $=\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})-\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},u_{t}))$ (19)
$\leq\langle\nabla_{w}\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t}),w_{t}-u_{t}\rangle$ (20)
$=\langle E[\hat{G}_{t}(w_{t})|S_{1:t-1}],w_{t}-u_{t}\rangle.$ (21)
Rest of the proof is identical to Bai et al. (2022).
Theorem 6 (Regret convergence for ATLAS-OFU).
Let $f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w)$ denote a network with the same feature extractor as that of $f_{t}^{\prime\prime}$ and a last linear layer with weight $w$. Let $f^{\rm atlas-ofu}=f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})$, where $w_{t}$ is the weight maintained at round $t$ by Algorithm 6. If we run Algorithm 1 with ATLAS in (Bai et al., 2022) and set up the step size pool =$\mathcal{H}=\{\eta_{i}=O\left(\frac{\sigma}{\sqrt{KT}}\right)\cdot 2^{i-1}|i\in[N]\}$ ($N=1+\lceil\frac{1}{2}\log_{2}(1+2T)\rceil$), then under the same assumptions as Lemma 1 in (Bai et al., 2022), UOGD-OFU satisfies that
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f^{\rm atlas-ofu};\mathcal{P}^{\rm test}_{t})-\frac{1}{T}\sum_{t=1}^{T}\min_{w\in\mathcal{W}}\ell(f(\cdot;\theta^{\rm feat}_{f_{t+1}^{\prime\prime}},w);\mathcal{P}^{\rm test}_{t})\right]\leq O\left(\left(\frac{K^{1/3}}{\sigma^{2/3}}+1\right)\frac{V_{T,\ell}^{1/3}}{T^{1/3}}+\sqrt{\frac{K}{\sigma^{2}T}}\right),$ (22)
where $V_{T,\ell}:=\sum_{t=2}^{T}\sup_{w\in\mathcal{W}}|\ell(f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w);\mathcal{P}_{t}^{\rm test})-\ell(f(\cdot;\theta^{\rm feat}_{f_{t-1}^{\prime\prime}},w);\mathcal{P}_{t-1}^{\rm test})|$, $\sigma$ denotes the minimum singular value of the invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$ and $K$ is the number of classes and the expectation is taken with respect to randomness in the revealed co-variates.
The proof is similar to that of Theorem 5 and hence omitted.
Discussion about the assumption.
In the theorems for UOGD and ATLAS, the definition of $V_{T,\ell}$ is shift severity from $\mathcal{P}^{\rm test}_{t}$. However, in the theorems for UOGD-OFU and ATLAS-OFU above, $V_{T,\ell}$ is shift severity from both $\mathcal{P}^{\rm test}_{t}$ and $\theta_{f_{t}^{\prime\prime}}^{\rm feat}$, which can be much larger. This might lead to harder convergence of the regret.
Appendix E Additional experiments
E.1 Additional details of datasets
Severity of CIFAR-10C in the experiment.
For each type of corruption in CIFAR-10C, we select a mild level and a high level of severity in the experiment section. Here we introduce the exact parameters of mild and high levels of severity for those corruptions. For Gaussian Noise, the severity levels for [mild, high] are [0.03, 0.07]. For Fog, the severity levels for [mild, high] are [(0.75,2.5), (1.5,1.75)]. For Pixelate, the severity levels for [mild, high] are [0.75, 0.65].
Details of additional datasets.
In addition to CIFAR-10 and CIFAR-10C, we evaluate on three more datasets: STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), and EuroSAT (Helber et al., 2019). Similar to CIFAR-10, we split the original train sets of these datasets into the train set and the validation set by the ratio $4:1$ and use the original test sets for sampling test images in the online test stage.
E.2 Results on Additional Datasets
Figure 2(b) and Figure 5 show the results for three additional datasets: STL10, CINIC, and EuroSAT, also with OLS-OFU (rotation degree prediction) under sinusoidal shift setting. In-line with CIFAR-10, we observe that OLS and OLS-OFU can perform better than Base and OFU, and OLS-OFU can outperform OLS. We find this pattern to be more consistent for EuroSAT and STL10 than CINIC.
Figure 5: Results on three more datasets. SSL in OLS-OFU is rotation degree prediction and online shift pattern is Bernoulli shift.
E.3 More Results on CIFAR-10
Figure 6: Results of Bernoulli shifts on CIFAR-10. OLS-OFU is evaluated with three SSL methods.
Figure 6 shows the results on CIFAR-10 for Bernoulli shift cross three SSL methods in OLS-OFU. Similar to Figure 2(a), OLS-OFU within all three SSL methods outperforms all baseline methods.
E.4 More Results on CIFAR-10C
Figure 7: Results of two online shift patterns on CIFAR-10C and three SSL methods in OLS-OFU.
We evaluate three SSL methods in OLS-OFU on CIFAR-10C for two online shift patterns. We pick moderately high severity levels for evaluating CIFAR-10C. In Figure 7 we can observe the consistent improvement from OLS to OLS-OFU and the SOTA performance of OLS-OFU.
E.5 More Results on CIFAR-10C with High Severity
Figure 8: Results of two online shift patterns on CIFAR-10C (high severity) and three SSL methods in OLS-OFU.
We evaluate three SSL methods in OLS-OFU on CIFAR-10C with high severity for two online shift patterns. In Figure 8, we can observe when SSL in OLS-OFU is rotation degree prediction or MoCo, the improvement from OLS to OLS-OFU is very significant but OLS-OFU cannot outperform OFU. The conclusion is similar to the discussion for Figure 4 in Section 4. However, OLS-OFU with SSL entropy minimization has different behavior. When the corruption of CIFAR-10C becomes more severe, OLS-OFU entropy minimization shows less improvement from OLS, if we compare the transition: Figure 7 (clean CIFAR-10), Figure 7 (CIFAR-10C with mild severity), Figure 8 (CIFAR-10C with high severity). This suggests that rotation degree prediction and MoCo are more appropriate SSL to address the domain shift from CIFAR-10 to CIFAR-10C.
E.6 Empirical evaluation for Equation 8
Figure 9 shows the comparison between LHS (OLS-OFU) and RHS (OLS) of the inequality in Equation 8 when the SSL in OLS-OFU is entropy minimization or MoCo. Similar to what we observe in Figure 3, the inequality in Equation 8 holds cross 4 data settings and two online shift patterns when OLS-OFU is implemented with entropy minimization or MoCo.
Figure 9: Sanity check for the inequality $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;{\color[rgb]{0.84,0.09,0.41}f_{t}^{\prime\prime}},q_{t}/q_{0});\mathcal{P}_{t}^{\rm test})\right]<\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{0},q_{t}/q_{0});\mathcal{P}_{t}^{\rm test})$ (Equation 8) when SSL in OLS-OFU is entropy minimization or MoCo.
E.7 Ablation for the order of updates and predictions
In the default framework of online distribution shift as shown in Figure 1, the model updates happen after the predictions for the samples at time step $t$. We would like to see if the model updated before the predictions would bring benefit. Figure 10 shows the comparison between ”predict first” and ”update first”. We can observe that there is no strong evidence to demonstrate the advantage of ”predict first” or ”update first” – the difference is indeed insignificant. However, because in the framework of ”predict first” OLS and OLS-OFU enjoy strong theoretical guarantees, we recommend ”predict first” in practice.
Figure 10: Ablation for the order of updates and predictions.
E.8 Details of SSL Methods
When the SSL loss is rotation degree prediction, it requires another network $f^{\rm deg}$ to predict the rotation degree, sharing the same feature extractor $\theta^{\rm feat}$ as $f_{0}$ but with a different set of downstream layers. Its SSL loss $\ell_{\rm ssl}(S;f)$ is defined as $\sum_{x\in S}\ell_{ce}(f^{\rm deg}(R(x,i)),i)$, where $i$ is an integer uniformly sampled from $[4]$, and $R(x,i)$ is to rotate $x$ with degree $\mathrm{DL}[i]$ from a list of degrees $\mathrm{DL}=[0,90,180,270]$. Alternatively, if the SSL loss is entropy minimization, $\ell_{\rm ssl}(S;f)$ would be the entropy $\sum_{x\in S}\sum_{k=1}^{K}f(x)_{k}\log f(x)_{k}$. Moreover, the SSL loss would be a contrastive loss (InfoNCE) where the positive example $x^{\prime}$ is an augmented version of $x$ and other samples in the same time step can be the negative examples. However, the batch size for a time step is small, e.g. $10$ in our experiment. MoCo updates with such a small batch won’t work. Thus, we experiment with MoCo by applying a batch accumulation strategy, which we introduce next.
E.8.1 Batch accumulation for MoCo
Prior test-time training (rotation degree prediction, entropy minimization) methods operate with the OFU framework easily, with feature extractor updates taking place in each time step. Rotation degree prediction originates as a self-supervised training method (Gidaris et al., 2018), thus naturally we evaluate more recent self-supervised training methods, namely MoCo (He et al., 2020; Chen et al., 2020b, 2021). No prior work shows how to use MoCo (or self-supervised learning in general) in a test-time training setting. Given that self-supervised training is sensitive to batch size, the intuition is that a larger batch size (much larger than the number of online samples available per time step) is required to perform a valid gradient update for a MoCo checkpoint. As such, we evaluated a batch accumulation strategy OLS-OFU (BA=$\tau$), where we continue evaluating online samples per time step, but only perform the online feature update at every $\tau$ steps (having accumulated the online samples throughout the $\tau$ steps in one batch). In particular, we perform feature extractor update every $\tau=50$ steps (for 1000 steps, feature update occurs 20 times, online samples evaluation occurs 1000 times), evaluating with 10 online samples per time step, using a smaller learning rate (0.0005) but test-time train with 10 epochs. Notice that $\tau=1$ is the default setting in Algorithm 1. OLS-OFU with MoCo presented in Figure 2, Figure 6 and Figure 7 is equivalent to OLS-OFU (BA=$50$).
To show the necessity of large $\tau$, we evaluate both OLS-OFU (BA=1) and OLS-OFU (BA=50) on CIFAR-10 (Figure 11) and CIFAR-10C (Figure 12). Firstly, we can observe that OLS-OFU (BA=1) is even worse than OLS. We hypothesize this is because small batch size of MoCo will hurt the performance and larger batch size in MoCo is necessary. Hence, we increase $\tau$ from $1$ to $50$ and then we can observe the significant improvement from OLS-OFU (BA=1) to OLS-OFU (BA=50). Now, OLS-OFU (BA=50) can outperform OLS.
Figure 11: Evaluating OLS-OFU with different batch accummulations for MoCo on CIFAR-10.
Figure 12: Evaluating OLS-OFU with different batch accummulations for MoCo on CIFAR-10C.
E.9 Improvements compared with literature
We would like to compare the improvements in the literature for the OLS problem and the improvements from the SOTA in the literature to our method OLS-OFU. We can calculate the difference $\Delta_{\rm error}$ between the errors of method A and B to show the improvement $A\to B$. To measure the improvements in the literature for the OLS problem, we measure $\Delta_{\rm error}$ between FTH (proposed in the first OLS paper) and FLHFTL (SOTA in the literature). We also calculate $\Delta_{\rm error}$ between FLHFTL and FLHFTL-OFU.
[…]
Figure 13: Improvements from FTH to FLHFTL and improvements from FLHFTL to FLHFTL-OFU across all datasets and online shift patterns.
E.10 Self-training
Pseudo labelling (Lee, 2013), a common self-training technique, generates pseudo labels for unlabelled data and uses them to update the model. Though we are not able to use ground-truth labels to compute feature extractor updates, we can use the model at time $t$ to make predictions with respect to the online samples at time $t$, and train on the inputs with their assigned (pseudo) labels. An issue that arises in self-training is confirmation bias, where the model repeatedly overfits to incorrect pseudo-labels. As such, different methods can be used to select which samples will be pseudo-labelled and used in updating the model, e.g. using data augmentation (Arazo et al., 2020), using regularization to induce confident low-entropy pseudo-labelling (Grandvalet & Bengio, 2004a), using softmax thresholds to filter out noisy low-confidence predictions (Xie et al., 2020). We make use of ensembles to identify noisy low-confidence/entropy pseudo-label predictions, though other various alternatives can also be used. In addition to OLS and OLS-OFU, we highlight the methods under comparison:
  • OLS-OFU ($\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$): Instead of computing pseudo-labels, we make use of the correct ground-truth labels $y_{\textnormal{ground-truth}}$. Recall $\ell_{\rm sup}$ is the supervised learning loss. We update the feature extractor with the supervised loss w.r.t. ground-truth labels $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$.
  • OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$): Instead of computing pseudo-labels, we make use of the correct ground-truth labels $y_{\textnormal{ground-truth}}$. Recall $\ell_{\rm ssl}$ and $\ell_{\rm sup}$ are the self-supervised and supervised learning losses respectively. We update the feature extractor with both the self-supervised loss $\ell_{\rm ssl}$ as well as the supervised loss w.r.t. ground-truth labels $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$.
  • OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{pseudo-label(\#samples=, \#FU-samples=)}})$): Recall $\ell_{\rm ssl}$ and $\ell_{\rm sup}$ are the self-supervised and supervised learning losses respectively. We compute pseudo-labels $y_{\textnormal{pseudo-label}})$, and update the feature extractor with both the self-supervised loss $\ell_{\rm ssl}$ as well as the supervised loss w.r.t. pseudo-labels $\ell_{\rm sup}(\cdot,y_{\textnormal{pseudo-label}})$.
How to compute pseudo-labels?
We now describe the procedure to compute pseudo-labels for $\ell_{\rm sup}(\cdot,y_{\textnormal{pseudo-label(\#samples=, \#FU-samples=)}})$. The seed used to train our model is 4242, and we train an additional 4 models on seeds 4343, 4545, 4646, 4747. With this ensemble of 5 models, we keep sampling inputs at each online time step until we have #FU-samples samples, or we reach a limit of #samples samples. We accept an input when the agreement between the ensembles exceeds a threshold $e=1.0$ (i.e. we only accept samples where all 5 ensembles agree on the label of the online sample). In the default online learning setting, there are only #samples=10, and therefore there may not be enough accepted samples to perform feature update with, thus we evaluate with a continuous sampling setup, where we sample #samples=50 (and evaluate on all these samples), but only use the first 10 samples (#FU-samples=10) to perform the feature extractor update.
Results on pseudo-labelling.
First, we find that OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$) attains the lowest error and is the lower bound we are attaining towards. Evaluating OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{pseudo-label(\#samples=10, \#FU-samples=10)}})$), we find that the performance does not outperform OLS-OFU, and is not near OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$). If we set the threshold $e$ too high, there may not be enough online samples to update the feature extractor. If we set the threshold $e$ too low, there may be too many incorrect labels and we incorrectly update our feature extractor. As such, we would like to sample more inputs at each online time step such that we can balance this tradeoff. We sample #samples=50 at each online time step, and update with #FU-samples $\leq$ 10. For fair comparison, we also show the comparable methods in both #samples=10, #FU-samples=10 and #samples=50, #FU-samples=50 settings.
With this sampling setup, we find that OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{pseudo-label(\#samples=50, \#FU-samples=10)}})$) can outperform both OLS-OFU (#samples=10) and OLS-OFU (#samples=50). Though it does not exceed neither OLS-OFU ($\ell_{\rm ssl}$ + $\ell_{\rm sup}(\cdot,y_{\textnormal{ground-truth}})$) for #samples=10 nor #samples=50, it lowers the gap considerably.
Figure 14: Results on pseudo-labelling.